A special novelty coin has a probability of 0.89 of coming up heads. In 12 tosses of this coin:

a) What is the probability the coin comes up heads exactly 10 times?

Round your response to at least 3 decimal places.

   

b) What is the probability the coin comes up heads more than 10 times?

Round your response to at least 3 decimal places.
Answer #1

This will be a binomial experiment with parameters:

n = 12, p = 0.89

a) P(Exactly 10)

= (12) (0.89)10 (0.11) 10

= 0.249

b) P(More than 10)

= P(11) + P(12)

= (12) (0.89)}(0.11) + (12) (0.89) 12 (0.11)

= 0.613